OpenStudy (anonymous):

Calculate the mass of a rectangular solid that has a density of 2.53 g/cm3 and which measures 25 mm by 1.80 cm by 3.00 cm. Dont say in what but i assume grams

OpenStudy (anonymous):

density = mass / volume \[volume = 0.25cm * 1.8 cm * 3.00cm\] this is the volume of the rectangular solid and beware of units density = 2.53 g /cm3 So 2.53 = m / (.25 * 1.8 * 3) m = 1.87 g hope this is what you are looking for.

OpenStudy (anonymous):

2.5 cm not .25 Right?

OpenStudy (anonymous):

yeah, 2.5 cm.

OpenStudy (anonymous):

Sorry, it is 2.5 cm

OpenStudy (unklerhaukus):

\[\frac mV=\rho\] \[m=\rho\times V \]\[=2.53 \left[\frac{\text g}{{\text {cm}}^3}\right] \times 2.5[\text {cm}]\times1.80 [\text {cm}]\times3.00 [\text {cm}]\]\[=(2.53\times2.5\times1.8\times3)[\text g]\]

OpenStudy (anonymous):

Thanks :) but you didn't have to do all that equation editing :P

OpenStudy (unklerhaukus):

did you get the right answer ?

OpenStudy (anonymous):

34.1550 g